Emergency Switch Replacement (Status)

« Back

[#1015] Emergency Switch Replacement (Status)

Posted: 2023-10-29 22:17

Start: 2023-10-29 22:00:00
End : 2023-10-29 23:20:00

Affects: Servers Rack B-2-13 NZS

We've found an issue with both optics leading to our rackswitch in Rack B-2-13, in Nedzone Steenbergen.

After replacing optics and fibers, as well as cleaning them, there are still errors occurring.

We're currently preparing a new switch to seamlessly move servers onto.

Please expect an update here once it is done.

Update 23:26 - Switch has been replaced and network traffic is restored. If you experience any issues, please contact us through Ticketing.